OpenStudy (anonymous):

Find the domain of y=2/3x-1 if the range is {-3, -1, 7, 17}.

OpenStudy (solomonzelman):

plug in -3,-1,7,17 for y, and find the x. For example. lets first rearrange the equation (for all range and domain). (by multiplying both sides time (3x-1). y(3x-1)=2 when the range is -3, (-3)(3x-1)=2 -9x+3=2 -9x=-1 x= 1/9 So you would start filling in the domain as, {1/9, ...... } you just need to find the rest of the values of the domain. When x=-1? When x=7? When x=17?

OpenStudy (texaschic101):

exactly..the domain is all the x values and the range is all the y values. So if you are given a range of -3, and the range is the y values, then sub -3 in for y and solve for x, the domain.